Calculate the volume of a sphere with a radius of 347pm in cubic meters


1
Expert's answer
2020-08-31T07:05:31-0400

Solution.

"R = 347 \ pm = 3.47\times 10^{-10} \ m"

"V = \frac{4 \times \pi \times R^3}{3}"

"V = 1.75 \times 10^{-82} \ m^3"

Answer:

"V = 1.75 \times 10^{-82} \ m^3"